OpenStudy (anonymous):

Determine the constants a and b, given that the parabola y = ax^2+bx+3 passes through (−2, 34)and (3,9). a=? b=?

OpenStudy (tkhunny):

Substitute the known values and solve the resulting system.

OpenStudy (ammarah):

Help me somepne please!!!

OpenStudy (saifoo.khan):

Insert eh values of x and y given into the equation one-by-one. You willl then get 2 equations. Solve both of them simultaneously.

OpenStudy (saifoo.khan):

Like this: y = ax^2+bx+3 passes through (−2, 34)and (3,9). 34 = a(-2)^2 + b(-2) + 3 and 9 = a(3)^2 + b(3) + 3 Simplify both.

OpenStudy (anonymous):

Thank You so much! :)